OpenStudy (anonymous):

Evaluate the series sin^2 0 + sin^2 1 + sin ^2 2+ sin^2 3... sin^2 90

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \sum_{ n=1 }^{ 90 } ~ {\rm Sin}^2({\rm n})}\)

OpenStudy (solomonzelman):

you can enter a series into wolfram and it will give you the sum, if you are allowed to do that.

OpenStudy (solomonzelman):

My interpretation is off, it should say \(\large\color{black}{ \displaystyle \sum_{ n=0 }^{ \infty } ~ {\rm Sin}^2({\rm n})}\)

OpenStudy (anonymous):

thanks @solomonzelman

OpenStudy (solomonzelman):

sure